# Enhancing Accuracy and Robustness of Steering Angle Prediction with Attention Mechanism

### 16 Dec 2023 v3

@@ V-B Qualitative
Figure 9: Saliency maps representing the focus areas of the different models. The saliency maps use three colors. Black signifies no focus, red indicates some focus, and yellow indicates the highest level of focus. To generate a saliency map we get a predicted output from the model. We take the max score of this output and perform back-propagation on the max score with respect to the original input. We use layer 4 saliency as an example. For the last three subfigures, the saliency maps are blended with the original image with a blending ratio of 0.75.
VI Conclusion and Future Work
This research first examines the two main veins of modern neural network architectures and their intra-variations. The models explored for this research include various InceptionNet and ResNet depths of complexity. If there is no limit on the model/data size, then it appears that the ResNet architecture tends to perform better with deeper and more complex structures. In addition, we have also added attention modules to ResNet architectures, which has further increased the models’ performance by decreasing the MSE up to 6.83% for the SAP dataset and up to 6.09% for the custom dataset. All of the programs and the custom dataset used in this study are publicly accessible in this repository [21]. Our investigation extended into the realm of adversarial robustness, where we systematically tested the resilience of our models against adversarial attacks, thereby uncovering potential vulnerabilities. Intriguingly, the incorporation of attention modules demonstrated a promising capability to fortify the models against adversarial attacks, thereby reducing their adverse impact on predictive accuracy. In our future work, we will explore different defense mechanisms to overcome these adversarial attacks.

### 01 Feb 2024 v4

@@ IV-C Qualitative Results
Figure 7: Saliency maps representing the focus areas of the model with and without our introduced attention modules. Black signifies no focus, red indicates some focus, and yellow indicates the highest level of focus. To generate a saliency map we get a predicted output from the model. We take the max score of this output and perform back-propagation on the max score with respect to the original input. We use layer 4 saliency as an example. For ResNet32 and ResNet32 with attention, the saliency maps are blended with the original image with a blending ratio of 0.75.
To better understand the influence of our introduced attention modules, we illustrated the changes in saliency maps, which provide insight into where the model is focused on in an image (Fig. 7). According to the results, we can see that the attention modules assist ResNet in directing its focus toward the road ahead, prioritizing it over irrelevant areas. Sometimes (Row 2 and Row 4 in Fig. 7), without our attention mechanism, the original ResNet model focuses only on one side of the road, and with the attention mechanism added, the model can pay attention to the center and the other side of the road to a greater extent. Fig. 7 qualitatively explained our approach’s efficacy.
V Conclusion
Targeting the task of steering angle prediction, this paper first examines the two main veins of modern neural network architectures and their intra-variations of various complexity (i.e., InceptionNets and ResNets). In addition, we have introduced attention modules to deep steering angle predictors, which has increased the models’ performance by decreasing the MSE up to 6.83% for the SAP dataset and up to 6.09% for the custom dataset. The incorporation of our attention modules also demonstrates a promising capability to boost steering angle predictors’ adversarial robustness. For example, it can enhance the adversarial robustness of ResNet32 by up to 53.96% against the PGD attack.